OpenStudy (anonymous):

A particle is moving with velocity v(t) = t^2 – 9t + 18 with distance, s measured in meters, left or right of zero, and t measured in seconds, with t between 0 and 8 seconds inclusive. The position at time t = 0 sec is 1 meter right of zero, that is, s(0) = 1. The average velocity over the interval 0 to 8 seconds The instantaneous velocity and speed at time 5 secs The time interval(s) when the particle is moving right The time interval(s) when the particle is going faster slowing down Find the total distance the particle has traveled between 0 and 8 seconds

OpenStudy (anonymous):

The instantaneous velocity and speed at time 5 secs is v(t)=\[t=5 , v(t)=t^2-9t+18 =(5)^29(5)+18= 25-45+18=/-2/=2\]

OpenStudy (irishboy123):

*The average velocity over the interval 0 to 8 seconds* work out how far it actually travels in 8 sec and divide by 8. integration: bearing in mind, "time t = 0 sec is 1 meter right of zero"

OpenStudy (anonymous):

v(t) = t^2 – 9t + 18 average velocity = (t^2-9t+18)/8=(8^2-9*8+18)/8 instantaneous velocity v(5)=5^2-9*5+18 you could plot the funtion V vs. t

OpenStudy (anonymous):

@AJ01 what do you mean V vs. t? i am sorry i have been out off school for a week don't remember anything from that class.

OpenStudy (freckles):

Like he wants you to graph v(t)=t^2-9t+18

OpenStudy (anonymous):

V: velocity vs. time

OpenStudy (perl):

since you are finding total distance, you want to integrate the absolute value of velocity

OpenStudy (perl):

integral |v(t) |

OpenStudy (anonymous):

@perl /V(t)/=2

OpenStudy (perl):

total distance should be 107/3

OpenStudy (perl):

35.67

OpenStudy (perl):

are you allowed to use a calculator?

OpenStudy (perl):

do you have a TI 83

OpenStudy (anonymous):

@perl yes, but i don't have it with me now.

OpenStudy (perl):

Math->9:fnInt( fnInt(abs(x^2-9*x+18, x , 0, 8 ) ) = 35.666666666... = 107/3

OpenStudy (perl):

when you get a chance take out your calculator

OpenStudy (anonymous):

@perl thank you soooooooo much.